OpenStudy (anonymous):

Medal & fan. The sum of the squares of 2 consecutive negative integers is 41. What are the numbers? Which of the following equations is the result of using the factoring method to solve the problem?

OpenStudy (andrearchat):

options?

OpenStudy (anonymous):

it would be 4.5*2

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

(n - 5)(n - 4) = 0 (n - 5)(n + 4) = 0 (n + 5)(n - 4) = 0 (n + 5)(n + 4) = 0

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

it can only be the first one because they are asking for two consecutive integers

OpenStudy (anonymous):

so it would be (n-5)(n-4)=0

OpenStudy (anonymous):

medals?

OpenStudy (anonymous):

thnx